# Regularization of the ensemble Kalman filter using a non-parametric, non-stationary spatial model

### v3

@@ 2.4 Positive definite convolution kernel
In E we consider the process $\xi_{\rm statio}$ defined by the process convolution model with the stationary kernel $u(\rho)$, see Eq. (46). As discussed above, the kernel is not unique given the covariance function of the process. In an attempt to select a unique kernel, we impose a computationally motivated spatial localization requirement, seeking $u(\rho)$ that has the smallest spatial scale. We show that there are three solutions to this optimization problem, all of them having the same shape of the kernel. One of these three equivalent solutions is positive definite, which we select. A similar result is valid for the process defined on the circle. So, with a stationary process defined by the convolution $\xi=u*\alpha$ (where $\alpha$ is the white noise) and having a given spatial covariance function, the most spatially localized kernel $u$ can be considered to be a positive definite function of distance.
@@ 2.5 Uniqueness of the convolution kernel
In F we prove that under the local isotropy constraint (section 2.3), the positive-definite-kernel constraint (section 2.4), and a technical assumption, the kernel $u(x,\rho)$ is uniquely determined by the model’s non-stationary covariances $B(x,x^{\prime})$.
@@ 2.6 Local stationarity
We say that a process is locally stationary if it can be approximated by a stationary process locally in space, see G for details. To formulate conditions under which the process defined above by Eq. (6) is locally stationary, we take advantage of the model formulation based on the locally isotropic convolution kernel $u(x,\rho(x,y))$, Eq. (3). That formulation enables us to introduce two spatial scales. The first one is the length scale at which $u(x,\rho)$ typically varies as a function of distance $\rho$. We call it the characteristic length scale $L$ of the process $\xi(x)$ (defined in G). The other one is the length scale at which $u(x,\rho)$ varies as a function of location $x$. We refer to it as the non-stationarity length scale $\Lambda$, also introduced in G. We show that the process $\xi(x)$ is locally stationary if one of the following two conditions are met.
The first (main) condition is $L\ll\Lambda$, that is, the requirement that the local structure of the process changes in space on a scale much larger than the length scale of the process itself. The second (less interesting and rather trivial) condition is that the spatial variability in the convolution kernel (measured by the non-stationarity strength, see G) is small. In this case, the process is locally stationary even if $\Lambda$ is comparable with $L$.
These conditions are obtained in G under the assumption that the local spectral functions $\sigma_{0}(x),\sigma_{1}(x),\dots,\sigma_{\ell_{\rm max}}(x)$ are jointly a stationary vector-valued random process of location $x$ and considering the conditional distribution of the process $\xi$ given its local spectra. Equivalently, conditioning on the convolution kernel $u(x,\rho)$ (viewed as a Hilbert space-valued stationary random process of $x$) can be considered.
@@
Local stationarity is the third constraint we introduce. Note that the slow variation of the kernel $u(x,\rho)$ with location $x$ (compared with its variation as a function of distance $\rho$), actually, justifies the term “local spectrum” (section 2.3) introduced first in the time series context by Priestley (1965), who called it evolutionary spectrum.
@@ 2.7 Smoothness of local spectra
Studies of real-world spatio-temporal processes showed that spatial (and temporal) spectra in turbulent flows are quite smooth, exhibiting typically a power-law behavior at large wavenumbers, e.g., Gage and Nastrom (1986), Trenberth and Solomon (1993). For this reason and to regularize the spatial model by further reducing its effective number of degrees of freedom, we postulate that the local spatial spectral functions $\sigma_{\ell}(x)$ are smooth functions of the wavenumber $\ell$ — this is our fourth constraint. As for a measure of smoothness of the local spectrum, we provide one definition in I and impose smoothness by placing a prior distribution over the functions $\sigma_{\ell}(x)$ in section 3.3.
Note that if $\sigma_{\ell}(x)$ is smooth as a function of $\ell$, then its inverse spectral transform, the convolution kernel $u(x,\rho)$, rapidly decreases with distance $\rho$. The rapidly decreasing convolution kernel prohibits spurious (and damaging in the analysis) long-distance correlations, which are present in sample (ensemble) covariances.
@@ 2.8 Summary of constraints
  4. 4.
    The local spatial spectra $f_{\ell}(x)$ are smooth functions of the wavenumber $\ell$.
@@ 3 Estimation of the spatial model from the ensemble
In this section, we propose an estimator of the Locally Stationary Convolution Model from an ensemble (sample) of $M$ independent identically distributed random field realizations (ensemble members): $\{\xi^{(\mu)}(x)\}_{\mu=1}^{M}$. The goal is to estimate the convolution kernel $u(x,\rho(x,\cdot))$, where $x$ runs over points of a spatial grid. In this study, we confine ourselves to point estimates of $u(x,\rho(x,y))$. Recall that for any $x$, the convolution kernel $u(x,\rho(x,y))$ is the inverse Fourier-Legendre transform of the spectral functions $\sigma_{\ell}(x)=\sqrt{f_{\ell}(x)}$ (see Eq. (4)), so it suffices to estimate the local spectra $f_{\ell}(x)$ or the spectral functions $\sigma_{\ell}(x)$. We seek an estimator potentially applicable to high-dimensional problems that have $10^{9}$ degrees of freedom or more.
We note that the exact likelihood of the data $\{\xi(x_{i})\}_{i=1}^{n_{\rm x}}$ (where $n_{\rm x}$ stands for the size of the spatial grid) generated by the model Eq. (6) is computationally intractable in high dimensions. Therefore, we rely on composite likelihood (Lindsay, 1988, Varin et al., 2011), arriving at a two-stage estimator introduced below in this section. First, several bandpass filters are applied to the input ensemble and sample variances of their outputs are computed. Second, the local spectra $f_{\ell}(x)$ are restored from those sample variances.
@@ 3.1 Multiple spatial bandpass filters
In the stationary case, it is straightforward to derive from Eq. (44) that the likelihood of the data, i.e., the ensemble of random-field realizations $\{\xi^{(\mu)}(x)\}_{\mu=1}^{M}$, given the spectrum $\{f_{\ell}\}_{\ell=0}^{\ell_{\rm max}}$, depends on the data through the set of single-wavenumber sample variances $\widehat{v}_{\ell}=\frac{1}{M}\sum_{\mu=1}^{M}\sum_{m=-\ell}^{\ell}|\widetilde{\xi}_{\ell m}^{(\mu)}|^{2}$. (Note that with $M=1$, the set of $\widehat{v}_{\ell}$ is the spherical analog of what is known in the time series literature as the periodogram.) Therefore, the vector of the single-wavenumber sample variances $\widehat{v}_{\ell}$ (with $\ell=0,1,\dots,\ell_{\rm max}$) is a sufficient statistic for the spectrum $f_{\ell}$. This implies that no information on the true spectrum is lost if we switch from the raw ensemble to the set of $\widehat{v}_{\ell}$.
@@
In the locally stationary case, however, relying on the single-wavenumber sample variances is not a good idea because isotropic spatial filters that isolate individual spectral components have non-local impulse response functions. Indeed, the filter’s response function is the backward Fourier-Legendre transform of its spectral transfer function $H(\ell)$. If $H(\ell)$ is equal to one at the single wavenumber $\ell_{0}$ and zero otherwise, then $h(\rho(x,y))=\frac{2\ell_{0}+1}{4\pi}P_{\ell_{0}}(cos(\rho(x,y)))$ (see Eq. (34)). Due to the non-locality of $P_{\ell_{0}}(cos(\rho))$, the filtered signal, which is the convolution of the input signal with $h(\rho(x,y))$, will contain significant contributions from over the whole global domain. This is not acceptable in the non-stationary case, in which we seek primarily local contributions.
Therefore, we replace $\ell_{\rm max}+1$ filters that isolate individual spectral components with a smaller number, $J$, of bandpass filters ${\cal H}_{j}$ ($j=1,2,\dots,J$) with localized impulse reasponse functions, $h_{j}(\rho)$, i.e., with broader spectral transfer functions, $H_{j}(\ell)$. In the estimator, we replace the raw data $\xi^{(\mu)}(x)$ with the set of the filtered fields $\varphi_{j}^{(\mu)}(x)=({\cal H}_{j}\xi^{(\mu)})(x)$.
@@
Broadening the filters’ transfer functions reduces the spectral resolution of the estimator. But, according to our fourth constraint (section 2.8), the spatial spectra are smooth, so high spectral resolution is not needed. For more details on the bandpass filters, see sections 3.2, 5, and 6.
3.2 Sample band variances
As discussed in the previous subsection, to obtain some spatial resolution in the estimation of $\sigma_{\ell}(x)$, we have to reduce the spectral resolution. We do so by relying on the outputs $\varphi_{j}(x)$ of several bandpass filters ${\cal H}_{j}$, with $j=1,2,\dots,J$, and replacing the likelihood of the full data $\xi(x)$ by the likelihood of the bandpass-filtered signals $\varphi_{j}(x)$. Here $\xi$ denotes one ensemble member.

### v5

@@ 3.4 Positive definite convolution kernel
In D, we consider the process $\xi_{\rm statio}$ defined by the process convolution model with the stationary kernel $u(\rho)$, see Eq. (54). As discussed above, the kernel is not unique given the covariance function of the process. In an attempt to select a unique kernel, we impose a computationally motivated spatial localization requirement, seeking $u(\rho)$ that has the smallest spatial scale. We show that there are three solutions to this optimization problem, all of them having the same shape of the kernel. One of these three equivalent solutions is positive definite, which we select. So, if a stationary process with a given covariance function is defined by the convolution $\xi=u*\alpha$ (where $\alpha$ is the white noise), then the most spatially localized kernel $u$ can be taken to be a positive definite function of distance.
@@ 3.5 Uniqueness of the convolution kernel
In E we prove that under the local isotropy constraint (section 3.3), the positive-definite-kernel constraint (section 3.4), and a technical assumption, the kernel $u(x,\rho)$ is uniquely determined by the model’s non-stationary covariance function $B(x,x^{\prime})$.
@@ 3.6 Local stationarity
The notion of local stationarity has been defined differently by different authors (most often for processes on the real line). The general idea is that a locally stationary process can be approximated by a stationary process locally, i.e., in the vicinity of any point in time (Mallat et al., 1998).
Starting from Dahlhaus (1997), the common approach to studying local stationarity is the use of the ‘infill’ asymptotics, in which the non-stationarity of the process $\xi(t)$ is determined by a function, $a(\tau)$, defined in rescaled time $\tau\in[0,1]$. Then, $\xi(t)$ is considered on $t=T\tau\in[0,T]$ as $T\to\infty$. The prototypical example is the autoregressive process $\xi_{t+1}=a(t/T)\,\xi_{t}+\eta_{t}$, where $t=0,1,\dots,T$ and $\eta_{t}\sim N(0,1)$. The mean time-scale of this process does not change as $T\to\infty$, whereas the non-stationarity pattern is infinitely stretched. As a result, for any point $\tau_{0}$ in rescaled time, $\xi_{t}$ can be approximated by a stationary process in the vicinity of the point $t=T\tau_{0}$, with the approximation error vanishing as $T\to\infty$.
On compact manifolds like the circle or the sphere, this rescaling cannot be used because of their compactness. Therefore, we propose a different approach to local stationarity with our process convolution model.
Recall that the non-stationary random process in question is defined by Eq. (10), where the non-stationarity comes from the dependencies of $\sigma_{\ell}$ on the location $x$. We assume that these are caused by $\sigma_{\ell}(x)$ being random processes themselves. More specifically, we postulate that $\sigma_{\ell}(x)$ are, jointly, a vector-valued mean-square differentiable stationary random process. Then, we obtain spatial non-stationarity by considering the conditional distribution of the process $\xi$ given all $\sigma_{\ell}(\cdot)$. We denote this conditioning by $\xi\,|\,\sigma$. In other words, we model spatial non-stationarity by assuming that $\boldsymbol{\sigma}(x)=\{\sigma_{\ell}(x)\}_{\ell=0}^{\ell_{\rm max}}$ is a random process and examine $\xi(x)=\sum\sigma_{\ell}(x)\,\widetilde{\alpha}_{\ell m}\,Y_{\ell m}(x)$ with a random realization of $\boldsymbol{\sigma}(x)$ being fixed. Equivalently, conditioning on the convolution kernel $u(x,\rho)$ (viewed as a Hilbert space-valued stationary random process of $x$) can be considered.
$\operatorname{\mathbb{E}}\left|\frac{\partial\sigma_{\ell}}{\partial x}\right|^{2}=\frac{{\mathsf{Var}}\,\sigma_{\ell}}{\Lambda_{\ell}^{2}}$ (12)
(where $\partial/\partial x$ is the spatial gradient operator) and the overall non-stationarity length scale $\Lambda$ from
$\sum_{\ell=-\ell_{\rm max}}^{\ell_{\rm max}}\frac{2\ell+1}{4\pi}\frac{{\mathsf{Var}}\,\sigma_{\ell}}{\Lambda_{\ell}^{2}}=\frac{\operatorname{\mathbb{E}}\xi^{2}}{\Lambda^{2}}.$ (13)
It is straightforward to prove that the process $\xi(x)$ is indeed locally stationary in the sense that for any point $x_{0}$, it can be approximated by a stationary process $\zeta(x;x_{0})$, with the mean-square approximation error vanishing as $\Lambda\to\infty$. To do so, we define $\zeta(x;x_{0})=\sum_{\ell}\sigma_{\ell}(x_{0})\sum_{m}\,\widetilde{\alpha}_{\ell m}\,Y_{\ell m}(x)$, that is, with the argument of $\sigma_{\ell}$ fixed at $x_{0}$. Then, the mean approximation error variance is
$\operatorname{\mathbb{E}}\operatorname{\mathbb{E}}[(\xi(x)-\zeta(x;x_{0}))^{2}\,|\,\sigma]=\sum_{\ell=-\ell_{\rm max}}^{\ell_{\rm max}}\frac{2\ell+1}{4\pi}\operatorname{\mathbb{E}}\left[\sigma_{\ell}(x)-\sigma_{\ell}(x_{0})\right]^{2}\leq\\ \rho^{2}(x,x_{0})\sum_{\ell=-\ell_{\rm max}}^{\ell_{\rm max}}\frac{2\ell+1}{4\pi}\operatorname{\mathbb{E}}\left|\frac{\partial\sigma_{\ell}}{\partial x}\right|^{2}=\frac{\rho^{2}(x,x_{0})}{\Lambda^{2}}\operatorname{\mathbb{E}}\xi^{2}.$ (14)
Here the last equality is due to Eqs. (12) and (13).
Theoretically, Eq. (14) shows that in the locally stationary limit, as $\Lambda\to\infty$ and $\operatorname{\mathbb{E}}\xi^{2}={\rm{const}}$, the approximation error vanishes globally, that is, for all points $x$ on the sphere. Practically, we wish a locally stationary process to have correlations that, approximately, depend only on the distance $\rho$ for distances within a process length scale, $L$ (whose definition is application dependent). Then, Eq. (14) implies that in applications, we may assume local stationarity if $L\ll\Lambda$.
@@
Local stationarity is the third constraint we introduce. Note that the slow variation of the kernel $u(x,\rho)$ with location $x$ (compared with its variation as a function of distance $\rho$), actually, justifies the term ‘local spectrum’ (section 3.3) introduced first in the time series context by Priestley (1965), who called it evolutionary spectrum.
@@ 3.7 Smoothness of local spectra
Studies of real-world spatio-temporal processes showed that spatial (and temporal) spectra in turbulent flows are quite smooth, exhibiting typically a power-law behavior at large wavenumbers (e.g. Gage and Nastrom, 1986, Trenberth and Solomon, 1993). For this reason and to regularize the spatial model by further reducing its effective number of degrees of freedom, we postulate that the spectral functions $\sigma_{\ell}(x)$ are smooth functions of the wavenumber $\ell$ — this is our fourth constraint. As for a measure of smoothness of the local spectrum, in section 4.4 we propose one definition aimed to facilitate the proof of consistency of the estimator, and in section 4.5 we impose another, more practically oriented smoothness constraint by placing a prior distribution over the functions $\sigma_{\ell}(x)$.
Note that if $\sigma_{\ell}(x)$ is smooth as a function of $\ell$, then its inverse spectral transform, that is, the convolution kernel $u(x,\rho)$, rapidly decreases with distance $\rho$. The rapidly decreasing convolution kernel prohibits spurious (and damaging in the analysis) long-distance correlations, which are present in sample (ensemble) covariances.
@@ 3.8 Summary of constraints
  4. (4)
    The spectral functions $\sigma_{\ell}(x)$ are smooth functions of the wavenumber $\ell$.
@@
3.9 Spatial discretization
When used in the analysis to specify the prior covariance matrix (see section 5), the model needs to be discretized, that is, defined on a spatial grid. This is done by approximating the convolution integral in Eq. (5) as a sum. In this proof-of-concept study, we restricted the band-limited field $\xi(\cdot)$ to the regular latitude-longitude grid described in section 6.3. In a real-world high-dimensional application, the key point will be to design a numerical scheme in which, for each grid point $x$, the sum runs over a relatively small number of neighboring grid points. This will make the resulting matrix operator (denoted in section 5 as $\widehat{\bf W}$) sparse. Developing such an algorithm will be done in a follow-up study using a multi-scale/multigrid approach. A non-ensemble version of the multi-scale technique (with static spatial covariances) has been successfully tested in an operational meteorological data assimilation system (Tsyrulnikov et al., 2019). In that application, we employed a non-regular, quasi-uniform grid on the sphere to maintain sparsity of the $\widehat{\bf W}$ matrix in the polar regions, where the latitude-longitude grid is dense.
@@ 4 Estimation of the spatial model from the ensemble
In this section, we propose an estimator of the Locally Stationary Convolution Model from an ensemble (sample) of $K$ independent identically distributed random field realizations (ensemble perturbations), $\{\xi^{(k)}(x)\}_{k=1}^{K}$. The goal is to estimate the convolution kernel $u(x,\rho(x,\cdot))$, where $x$ runs over points of the spatial grid. In this study, we confine ourselves to point estimates of $u(x,\rho(x,y))$. We recall that for any $x$, the convolution kernel $u(x,\rho(x,y))$ is the inverse Fourier-Legendre transform of the spectral functions $\sigma_{\ell}(x)=\sqrt{f_{\ell}(x)}$, see Eq. (6), so it suffices to estimate $\sigma_{\ell}(x)$. We seek an estimator that is applicable to high-dimensional problems with $10^{9}$ degrees of freedom or more.
The outline of this section is as follows. We start by motivating, in section 4.1, the application of a multi-scale bandpass filter to ensemble perturbations with the intention to extract aggregated local spectra. Then we explain how the (computationally intractable) likelihood of the raw data $\{\xi^{(k)}(x_{i})\}$ (with $i=1,2,\dots,n_{\rm x}$ and $k=1,2,\dots,K$) can be approximated by a much simpler composite likelihood (Lindsay, 1988, Reid, 2013), see section 4.2. After that, we propose a two-stage estimator, section 4.3. At the first stage, aggregated local spectra are computed as sample variances of the bandpass filtered fields. At the second stage, the local spectra are disaggregated from these sample variances for each grid point independently. We explore two approaches to the disaggregation. The first one, based on a linear estimating equation, allows us to prove consistency of the resulting model estimator, section 4.4. The second one relies on a neural network to estimate the local spectra, section 4.5. Finally, the spatially varying convolution kernel is computed from the estimated local spectra (by substituting them into Eq. (6)) and used in the analysis. The estimator’s workflow is summarized in section 4.6.
@@ 4.1 Multiple spatial bandpass filters
In the stationary case, it is straightforward to derive from Eq. (52) that the likelihood of the data, i.e., the ensemble of random-field realizations $\{\xi^{(k)}(x)\}_{k=1}^{K}$, given the spectrum $\{f_{\ell}\}_{\ell=0}^{\ell_{\rm max}}$, depends on the data through the set of spectral-space single-wavenumber sample variances $\widehat{v}_{\ell}=\frac{1}{K}\sum_{k=1}^{K}\sum_{m=-\ell}^{\ell}|\widetilde{\xi}_{\ell m}^{(k)}|^{2}$. (Note that with $K=1$, the set of $\widehat{v}_{\ell}$ is the spherical analog of what is known in the time series literature as the periodogram.) Therefore, the vector of the single-wavenumber sample variances $\widehat{v}_{\ell}$ (with $\ell=0,1,\dots,\ell_{\rm max}$) is a sufficient statistic for the spectrum $f_{\ell}$. This implies that no information on the true spectrum is lost if we switch from the raw ensemble to the set of $\widehat{v}_{\ell}$.
@@
In the locally stationary case, however, relying on the single-wavenumber sample variances is not a good idea because isotropic spatial filters that isolate individual spectral components have non-local impulse response functions. Indeed, the filter’s response function is the backward Fourier-Legendre transform of its spectral transfer function $H(\ell)$. If $H(\ell)$ is equal to one at the single wavenumber $\ell_{0}$ and zero otherwise, then $h(\rho(x,y))=\frac{2\ell_{0}+1}{4\pi}P_{\ell_{0}}(cos(\rho(x,y)))$ (see Eq. (6)). Due to the non-locality of $P_{\ell_{0}}(cos(\rho))$, the filtered signal, which is the convolution of the input signal with $h(\rho(x,y))$, will contain significant contributions from the whole global domain. This is not acceptable in the non-stationary case, in which we seek primarily local contributions.
Therefore, instead of $\ell_{\rm max}+1$ filters that isolate individual spectral components, we introduce a smaller number, $J$, of bandpass filters ${\cal H}_{j}$ ($j=1,2,\dots,J$) that have broader spectral transfer functions $H_{j}(\ell)$. Broader transfer functions imply spatially localized impulse response functions, $h_{j}(\rho)$, as required with a locally stationary process. Different $H_{j}(\ell)$ represent partly overlapping spectral bands that, jointly, cover the whole spectral range. Importantly, the filtered signals $\varphi_{j}(x)={\cal H}_{j}\xi(x)$ evaluated at a single grid point $x$ contain spatial information about the signal in question $\xi(x)$ in a vicinity of $x$. So, we can infer local spatial statistics of $\xi(\cdot)$ from the set of $\varphi_{j}(x)$ evaluated at a single point $x$.
On the other hand, the likelihood of $\xi(x)$ is computationally intractable because it depends on a huge number of parameters, that is, the spectral functions $\sigma_{\ell}(x)$ evaluated at all spatial grid points $x$. In the estimator, we propose to replace the raw data $\xi(x)$ with the set of filtered fields $\varphi_{j}(x)$ and, crucially, switch from the likelihood of the whole spatial fields to the likelihood of $\boldsymbol{\varphi}(x)=\{\varphi_{j}\}_{j=1}^{J}(x)$ evaluated at each grid point $x$ individually. As we show below in section 4.2, the likelihood of $\boldsymbol{\varphi}(x)$ depends, approximately, on $\sigma_{\ell}(x)$ evaluated at the same point. This reduces the global estimation problem to a series of local problems of much smaller dimensionality.
@@
We note that broadening the filters’ transfer functions reduces the spectral resolution of the estimator. But, according to our fourth constraint (section 3.8), the spatial spectra are smooth, so high spectral resolution is not needed. For more details on the bandpass filters, see sections 4.2, 6, and 7.
4.2 Approximate likelihood of the filtered signals